Sally and Bill work at different jobs. Sally earns $12 per hour and Bill earns $10 per hour. They each earn the same amount per week, but Bill works 3 more hours. How many hours a week does Bill work?
Answer by checkley77(12844)

12X=10(X+3)
12X=10X+30
12X-10X=30
2X=30
X=30/2
X=15 HOURS FOR SALLY.
15+5=18 HOURS FOR BILL.
PROOF:
12*15=10(15+3)
180=10*18
180=180